# Foundation Inference Models for Markov Jump Processes

### v2

@@ Abstract
Markov jump processes are continuous-time stochastic processes which describe dynamical systems evolving in discrete state spaces. These processes find wide application in the natural sciences and machine learning, but their inference is known to be far from trivial. In this work we introduce a methodology for zero-shot inference of Markov jump processes (MJPs), on bounded state spaces, from noisy and sparse observations, which consists of two components. First, a broad probability distribution over families of MJPs, as well as over possible observation times and noise mechanisms, with which we simulate a synthetic dataset of hidden MJPs and their noisy observation process. Second, a neural network model that processes subsets of the simulated observations, and that is trained to output the initial condition and rate matrix of the target MJP in a supervised way. We empirically demonstrate that one and the same (pretrained) model can infer, in a zero-shot fashion, hidden MJPs evolving in state spaces of different dimensionalities. Specifically, we infer MJPs which describe (i) discrete flashing ratchet systems, which are a type of Brownian motors, and the conformational dynamics in (ii) molecular simulations, (iii) experimental ion channel data and (iv) simple protein folding models. What is more, we show that our model performs on par with state-of-the-art models which are finetuned to the target datasets.
Our pretrained model is available online.
@@ 1 Introduction
Equation 1 is the so-called master equation of the MJP whose solutions are completely characterized by an initial condition $p_{\text{\tiny MJP}}(x,t=0)$ and the transition probability rates ($f:\mathcal{X}\times\mathcal{X}\rightarrow\mathbb{R}^{+}$).
@@
With these preliminaries in mind, we shall say that to infer an MJP from a set of (noisy) observations $\mathbf{z}(\tau_{1}),\dots,\mathbf{z}(\tau_{l})$ on the process $\mathbf{z}(t)$, recorded at some observation times $\tau_{1},\dots,\tau_{l}$, means to infer both the transition probability rates and the initial condition determining the hidden MJP $X(t)$ that best explains the observations. In practice, statisticians typically assume that they directly observe the coarse-grained process $X(t)$. That is, they assume they have access to the (possibly noisy) values $x_{1},\dots,x_{l}$, taken by $X(t)$ at the observation times $\tau_{1},\dots,\tau_{l}$ (see Section 2). We shall start from the same assumptions. Statisticians then tackle the inference problem by (i) defining some (typically complex) model that encodes, in one way or the other, equation 1 above; (ii) parameterizing the model with some trainable parameter set $\theta$; (iii) and updating $\theta$ to fit the empirical dataset.
One issue with this approach is that it turns the inference of hidden MJPs into an instance of an unsupervised learning problem, which, as history shows, is far from trivial (see Section 2). Another major issue is that, if one happens to succeed in training said model, the trained parameter set $\theta^{*}$ will usually be overly specific to the training set $\{(x_{1},\tau_{1}),\dots,(x_{l},\tau_{l})\}$, which means it will likely struggle to handle a second empirical process, even if the latter can be described by a similar MJP.
@@
Figure 1 contains snapshots from two empirical processes of very different nature. The figure on the left shows a set of observations (blue circles) recorded from the discrete flashing ratchet process (black line). The figure on the right shows the ion flow across a cell membrane, which jumps between different activity levels (blue line). Despite the vast differences between the physical mechanisms underlying each of these processes, the coarse-grained representations of the second one (black line) is abstract enough to be strikingly similar to the first one. Now, we expect that — at this level of representation — one could train a single inference model to fit each processes (separately). Unfortunately, we also expect that an inference model trained to fit only one of these (coarse-grained) processes, will have a hard time describing the second one.
@@
In this paper we will argue that the notion of an MJP description (in coarse-grained space) is simple enough, that it can be encoded into the weights of a single neural network model. Indeed, instead of training, in an unsupervised manner, a complex model which (somehow) encodes the master equation, on a single empirical process, we will train, in a supervised manner, a simple neural network model on a synthetic dataset that is composed of many different MJPs and hence implicitly encodes the master equation. We recently showed that such a supervised learning approach can be used to train a neural model to perform zero-shot inference of ODE systems  (Seifner et al., 2024). Below we demonstrate that it can also be used to train an MJP inference model, which we call Foundation Inference Model (FIM), to perform zero-shot inference of MJPs from empirical processes of very different kinds, which take values in state spaces of different sizes.
@@
In what follows, we first review both classical and recent solutions to the MJP inference problem in Section 2. We then introduce the FIM methodology in Section 3, which consists of a synthetic data generation process and a neural inference model. In Section 4 we empirically demonstrate that our methodology is able to infer MJPs from a discrete flashing ratchet process, as well as from molecular dynamics simulations and experimental ion channel data, all in a zero-shot fashion, while performing on par with state-of-the-art models which are finetuned to the target datasets. Finally, Section 5 closes the paper with some concluding remarks about future work, while Section 6 comments on the main limitations of our methodology.
@@
Figure 2: Foundation Inference Model (FIM) for MJP. Left: Graphical model of the FIM (synthetic) data generation mechanism. Filled (empty) circles represent observed (unobserved) random variables. The light-blue rectangle represents the continuous-time MJP trajectory, which is observed discretely in time. See main text for details regarding notation. Right: Inference model. The network $\psi_{1}$ is called $K$ times to process $K$ different time series. Their outputs is first processed by the attention network $\Omega_{1}$ and then by the FNN $\phi_{1}$, $\phi_{2}$ and $\phi_{3}$ to obtain the estimates $\mathbf{\hat{F}}$, $\log\text{Var}\,\mathbf{\hat{F}}$ and $\hat{\pi}_{0}$, respectively.
@@ 2 Related Work
The inference of MJP from noisy and sparse observations (in coarse-grained space) is by now a classical problem in machine learning. There are three main lines of research. The first (and earliest) one attempts to directly optimize the MJP transition probability rates, to maximize the likelihood of the discretely observed MJP via expectation maximization (Asmussen et al., 1996; Bladt and Sørensen, 2005; Metzner et al., 2007). Thus, these works encode the MJP inductive bias directly into their architecture. The second line of research leverages a Bayesian framework to infer the posterior distribution over the transition rates, through various Markov chain Monte Carlo (MCMC) algorithms (Boys et al., 2008; Fearnhead and Sherlock, 2006; Rao and Teg, 2013; Hajiaghayi et al., 2014). Accordingly, these simulation-based approaches encode the MJP inductive bias directly into their trainable sampling distributions. The third one, also Bayesian in character, involves variational inference. Within it, one finds again MCMC (Zhang et al., 2017), as well as expectation maximization (Opper and Sanguinetti, 2007) and moment-based (Wildner and Koeppl, 2019) approaches. More recently, Seifner and Sánchez (2023) used neural variational inference (Kingma and Welling, 2013) and neural ODEs (Chen et al., 2018) to infer an implicit distribution over the MJP transition probability rates. All these variational methods encode the MJP inductive bias into their training objective and, in some cases, into their architecture too.
Besides the model of Seifner and Sánchez (2023), which automatically infers the coarse-grained representation $X(t)$ from $D$-dimensional signals, all the solutions above tackle the MJP inference problem directly in coarse-grained space. However, we will also investigate below the conformational dynamics of physical systems for which the recorded data lies in a continuous space. To approach such type of problems, we will first need to define a coarse-grained representation of the state space of interest. Fortunately for us, there is a large body of works, within the molecular simulation community, precisely dealing with different methods to obtain such representations, and we refer the reader to e.gNoé et al. (2020) for a review. McGibbon and Pande (2015), for example, leveraged one such method to infer the MJP transition rates describing a molecular dynamics simulation via maximum likelihood. Alternatively, researchers have also treated the conformational states in these systems as core sets, and inferred phenomenological MJP rates from them (Schütte et al., 2011), or modelled the fast intra-state events as diffusion processes, indexed by a hidden MJP, and inferred the latter either via MCMC (Kilic et al., 2021; Köhs et al., 2022) or variational  (Horenko et al., 2006; Köhs et al., 2021) methods.
@@ 3 Foundation Inference Models
Our methodology has two components. First, a data generation process that encodes our believes about the class of realizable MJP we aim to model. Second, a neural network model that maps subsets of the simulated MJP observations onto the initial condition and rate matrix of their target MJPs. We will explore the details of these two components in the following sections.
3.1 Synthetic Data Generation Process
@@
Let us denote the size of the largest state space we include in our ensemble with $C$, and arrange all transition rates, for every MJPs within the ensemble, into $C\times C$ rate matrices. Let us label these matrices with $\mathbf{F}$. We define the probability of recording the noisy sequence $x_{1}^{\prime},\dots,x_{l}^{\prime}$, at the observation times $\tau_{1},\dots,\tau_{l}$, as follows
@@
$\prod\limits_{i=1}^{l}p_{\text{\tiny noise}}(x_{i}^{\prime}|x_{i},\rho_{y})p_{\text{\tiny MJP}}(x_{i}|\tau_{i},\mathbf{F},\pi_{0})p_{\text{\tiny grid}}(\tau_{1},\dots,\tau_{l}|\rho_{\tau})p_{\text{\tiny rates}}(\mathbf{F}|\mathbf{A},\rho_{f})p(\mathbf{A},\rho_{f})p(\pi_{0}|\rho_{0}).$ (2)
Next, we specify the different components of Eq. 2.
Distribution over rate matrices. The distribution $p_{\text{\tiny rates}}(\mathbf{F}|\mathbf{A},\rho_{f})$ over the rate matrices encodes our beliefs about the class of MJPs we expect to find in practice. We define it to cover MJPs with state spaces whose sizes range from 2 until $C$, because we want our FIM to be able to handle processes taking values in all those spaces. The distribution is conditioned on the adjacency matrix $\mathbf{A}$, which encodes only connected state spaces (i.e. irreducible embedded Markov chains only), and a hyperparameter $\rho_{f}$ which encodes the range of rate values within the ensemble. Specifically, we define the transition rates as $F_{ij}=a_{ij}f_{ij}$, where $a_{ij}$ is the corresponding entry of $\mathbf{A}$ and $f_{ij}$ is sampled from a Beta distribution with hyperparameter $\rho_{f}$. Note that these choices restrict the values of the transition rates within the ensemble to the interval $(0,1)$ and hence, they restrict the number of resolvable transitions within the time horizon. We refer the reader to Appendix B, where we specify the prior $p(\mathbf{A},\rho_{f})=p(\mathbf{A})p(\rho_{f})$ and its consequences, as well as give details about the sampling procedure.
@@
Distribution over initial conditions. The distribution $p(\pi_{0}|\rho_{0})$, with hyperparameter $\rho_{0}$, is defined over the $C$-simplex, and encodes our beliefs about the initial state (i.e. the preparation) of the system. It enters the master equation as the class probabilities over the states, at the start of the process, that is: $p_{\text{\tiny MJP}}(x,t=0)=\pi_{0}$. Appendix B provides the specifics.
@@
Distribution over observation grids. The distribution $p_{\text{\tiny grid}}(\tau_{1},\dots,\tau_{l}|\rho_{\tau})$, with hyperparameter $\rho_{\tau}$, gives the probability of observing the MJP at the times $\tau_{1},\dots,\tau_{l}$, and thus encodes our uncertainty about the recording process. Given that we do not know a priori whether the data will be recorded regularly or irregularly in time, nor we know its recording frequency, we define this distribution to cover the observation grids we expect to encounter in practice. Note that the number of observation points on the grid is variable. Please see Appendix B for details.
@@
Distribution over noise process. Just as the (instantaneous) solution of the master equation $p_{\text{\tiny MJP}}(x|t,\mathbf{F},\pi_{0})$, the noise distribution $p_{\text{\tiny noise}}(x^{\prime}|x,\rho_{y})$, with hyperparameter $\rho_{y}$, is defined over the set of metastable states $\mathcal{X}$. Recall that FIM solves the MJP inference problem directly in the coarse-grained space. The noise distributions then encodes both, possible measurement errors that propagate through the coarse-grained representation, or noise in the coarse-grained representation itself. We provide details of its implementation in Appendix B.
@@
We use the generative model, Eq. 2 above, to generate $N$ MJPs, taking values on state spaces with sizes ranging from 2 to $C$. We then sample $K$ paths per MJP, with probability $p(K)$, on the interval $[0,T]$, where $T$ is the observation-time horizon. The $j$th instance of the dataset thus consists of $K$ paths and is given by
$\mathbf{F}_{j}\sim p_{\text{\tiny rates}}(\mathbf{F}|\mathbf{A}_{j},\rho_{fj}),\,\,\text{and}\,\,\,\pi_{0j}\sim p(\pi_{0}|\rho_{0}),\,\,\text{with}\,\,(\mathbf{A}_{j},\rho_{fj})\sim p(\mathbf{A},\rho_{f}),$
$\text{so that}\quad\Big{\{}X_{jk}(t)\Big{\}}_{k=1}^{K}\sim\text{Gillespie}(\mathbf{F}_{j},\pi_{0j}),$ (3)
@@
3.2 Supervised Inference Models
In this subsection we introduce a neural network model that processes a set of $K$ time series of the form $\{(x_{k1}^{\prime},\tau_{k1}),\dots,(x_{kl}^{\prime},\tau_{kl})\}_{k=1}^{K}$, as generated by the process in Eq. 3 above, and estimates the intensity rate matrix $\mathbf{F}$ and initial distribution $\pi_{0}$ of the hidden MJP.

### v3

@@ Abstract
Markov jump processes are continuous-time stochastic processes which describe dynamical systems evolving in discrete state spaces. These processes find wide application in the natural sciences and machine learning, but their inference is known to be far from trivial. In this work we introduce a methodology for zero-shot inference of Markov jump processes (MJPs), on bounded state spaces, from noisy and sparse observations, which consists of two components. First, a broad probability distribution over families of MJPs, as well as over possible observation times and noise mechanisms, with which we simulate a synthetic dataset of hidden MJPs and their noisy observations. Second, a neural recognition model that processes subsets of the simulated observations, and that is trained to output the initial condition and rate matrix of the target MJP in a supervised way. We empirically demonstrate that one and the same (pretrained) recognition model can infer, in a zero-shot fashion, hidden MJPs evolving in state spaces of different dimensionalities. Specifically, we infer MJPs which describe (i) discrete flashing ratchet systems, which are a type of Brownian motors, and the conformational dynamics in (ii) molecular simulations, (iii) experimental ion channel data and (iv) simple protein folding models. What is more, we show that our model performs on par with state-of-the-art models which are trained on the target datasets.
Our pretrained model, repository and tutorials are available online [^1] https://fim4science.github.io/OpenFIM/intro.html.
@@ 1 Introduction
Equation 1 is the so-called master equation of the MJP whose solutions are completely characterized by an initial condition $p_{\text{\tiny MJP}}(x,t=0)$ and the transition rates $f:\mathcal{X}\times\mathcal{X}\rightarrow\mathbb{R}^{+}$.
@@
With these preliminaries in mind, we shall say that to infer an MJP from a set of (noisy) observations $\mathbf{z}(\tau_{1}),\dots,\mathbf{z}(\tau_{l})$ on the process $\mathbf{z}(t)$, recorded at some observation times $\tau_{1},\dots,\tau_{l}$, means to infer both the transition rates and the initial condition determining the hidden MJP $X(t)$ that best explains the observations. In practice, statisticians typically assume that they directly observe the coarse-grained process $X(t)$. That is, they assume they have access to the (possibly noisy) values $x_{1},\dots,x_{l}$, taken by $X(t)$ at the observation times $\tau_{1},\dots,\tau_{l}$ (see Section 2). We shall start from the same assumptions. Statisticians then tackle the inference problem by (i) defining some (typically complex) model that encodes, in one way or the other, equation 1 above; (ii) parameterizing the model with some trainable parameter set $\theta$; and (iii) updating $\theta$ to fit the empirical dataset.
@@
One issue with this approach is that it turns the inference of hidden MJPs into an instance of an unsupervised learning problem, which, as history shows, is far from trivial (see Section 2). Another major issue is that, if one happens to succeed in training said model, the trained parameter set $\theta^{*}$ will usually be overly specific to the training set $\{(x_{1},\tau_{1}),\dots,(x_{l},\tau_{l})\}$, which means it will likely struggle to handle a second empirical process, even if the latter can be described by a similar MJP. Figure 1 contains snapshots from two empirical processes of very different nature. The figure on the left shows a set of observations (blue circles) recorded from the discrete flashing ratchet process (black line). The figure on the right shows the ion flow across a cell membrane, which jumps between different activity levels (blue line). Despite the vast differences between the physical mechanisms underlying each of these processes, the coarse-grained representations of the second one (black line) is abstract enough to be strikingly similar to the first one. Now, we expect that — at this level of representation — one could train a single inference model to fit each process (separately). Unfortunately, we also expect that an inference model trained to fit only one of these (coarse-grained) processes, will have a hard time describing the second one.
@@
In this paper we will argue that the notion of an MJP description (in coarse-grained space) is simple enough, that it can be encoded into the weights of a single neural network model. Indeed, instead of training, in an unsupervised manner, a complex model (which somehow encodes the master equation) on a single empirical process; we will train, in a supervised manner, a simple neural network model on a synthetic dataset that is composed of many different MJPs, and hence implicitly encodes the master equation. This procedure can be understood as an amortization of the probabilistic inference process through a single recognition model, and is therefore akin to the works of Stuhlmüller et al. (2013)Heess et al. (2013) and Paige and Wood (2016). Rather than treating, as these previous works do, our (pretrained) recognition model as auxiliary to Monte Carlo or expectation propagation methods, we employ it to directly infer hidden MJPs from various synthetic, simulation and experimental datasets, without any parameter fine-tuning. We thus adopt the “zero-shot” terminology introduced by Larochelle et al. (2008), by which we mean that our procedure aims to recognize objects (i.e. MJPs) whose instances (i.e. noisy and sparse series of observations on them) may have not been seen during training. We have recently shown that such an amortization can be used to train a recognition model to perform zero-shot imputation of time series data (Seifner et al., 2024). Below we demonstrate that it can also be used to train a model of minimal inductive biases, to perform zero-shot inference of hidden MJPs from empirical processes of very different kinds, which take values in state spaces of different sizes. We shall call this recognition model Foundation Inference Model [^2] We name our model foundation model because it aligns with the definition proposed by Bommasani et al. (2021). Indeed, they define foundation models as any model that is trained on broad data (generally using self-supervision at scale) that can be adapted to a wide range of downstream tasks. (FIM) for Markov jump processes.
@@
In what follows, we first review both classical and recent solutions to the MJP inference problem in Section 2. We then introduce the FIM methodology in Section 3, which consists of a synthetic data generation model and a neural recognition model. In Section 4 we empirically demonstrate that our methodology is able to infer MJPs from a discrete flashing ratchet process, as well as from molecular dynamics simulations and experimental ion channel data, all in a zero-shot fashion, while performing on par with state-of-the-art models which are trained on the target datasets. Finally, Section 5 closes the paper with some concluding remarks about future work, while Section 6 comments on the main limitations of our methodology.
@@
Figure 2: Foundation Inference Model (FIM) for MJP. Left: Graphical model of the FIM (synthetic) data generation mechanism. Filled (empty) circles represent observed (unobserved) random variables. The light-blue rectangle represents the continuous-time MJP trajectory, which is observed discretely in time. See main text for details regarding notation. Right: Inference model. The network $\psi_{1}$ is called $K$ times to process $K$ different time series. Their outputs is first processed by the attention network $\Omega_{1}$ and then by the FNNs $\phi_{1}$, $\phi_{2}$ and $\phi_{3}$ to obtain the estimates $\mathbf{\hat{F}}$, $\log\text{Var}\,\mathbf{\hat{F}}$ and $\boldsymbol{\hat{\pi}}_{0}$, respectively.
@@ 2 Related Work
The inference of MJP from noisy and sparse observations (in coarse-grained space) is by now a classical problem in machine learning. There are three main lines of research. The first (and earliest) one attempts to directly optimize the MJP transition rates, to maximize the likelihood of the discretely observed MJP via expectation maximization (Asmussen et al., 1996; Bladt and Sørensen, 2005; Metzner et al., 2007). Thus, these works encode the MJP inductive bias directly into their architecture. The second line of research leverages a Bayesian framework to infer the posterior distribution over the transition rates, through various Markov chain Monte Carlo (MCMC) algorithms (Boys et al., 2008; Fearnhead and Sherlock, 2006; Rao and Teg, 2013; Hajiaghayi et al., 2014). Accordingly, these simulation-based approaches encode the MJP inductive bias directly into their trainable sampling distributions. The third one, also Bayesian in character, involves variational inference. Within it, one finds again MCMC (Zhang et al., 2017), as well as expectation maximization (Opper and Sanguinetti, 2007) and moment-based (Wildner and Koeppl, 2019) approaches. More recently, Seifner and Sánchez (2023) used neural variational inference (Kingma and Welling, 2013) and neural ODEs (Chen et al., 2018) to infer an implicit distribution over the MJP transition rates. All these variational methods encode the MJP inductive bias into their training objective and, in some cases, into their architecture too.
Besides the model of Seifner and Sánchez (2023), which automatically infers the coarse-grained representation $X(t)$ from $D$-dimensional, countinuous signals, all the solutions above tackle the MJP inference problem directly in coarse-grained space. Yet below, we also investigate the conformational dynamics of physical systems for which the recorded data lies in a continuous space. To approach such type of problems, we will first need to define a coarse-grained representation of the state space of interest. Fortunately for us, there is a large body of works, within the molecular simulation community, precisely dealing with different methods to obtain such representations, and we refer the reader to e.gNoé et al. (2020) for a review. McGibbon and Pande (2015), for example, leveraged one such method to infer the MJP transition rates describing a molecular dynamics simulation via maximum likelihood. Alternatively, researchers have also treated the conformational states in these systems as core sets, and inferred phenomenological MJP rates from them (Schütte et al., 2011), or modelled the fast intra-state events as diffusion processes, indexed by a hidden MJP, and inferred the latter either via MCMC (Kilic et al., 2021; Köhs et al., 2022) or variational  (Horenko et al., 2006; Köhs et al., 2021) methods.
@@ 3 Foundation Inference Models
Our methodology has two components. First, a data generation model that encodes our believes about the class of realizable MJPs we aim to model. Second, a neural recognition model that maps subsets of the simulated MJP observations onto the initial condition and rate matrix of their target MJPs. We will explore the details of these two components in the following sections.
3.1 Synthetic Data Generation Model
@@
Let us denote the size of the largest state space we include in our ensemble with $C$, and arrange all transition rates, for every MJPs within the ensemble, into $C\times C$ rate matrices. Let us label these matrices with $\mathbf{F}$. We define the probability of recording the noisy sequence $x_{1}^{\prime},\dots,x_{l}^{\prime}\in\mathcal{X}$, at the observation times $0<\tau_{1}<\dots<\tau_{l}<T$, with $T$ the observation time horizon, as follows
@@
$\prod\limits_{i=1}^{l}p_{\text{\tiny noise}}(x_{i}^{\prime}|x_{i},\rho_{x})p_{\text{\tiny MJP}}(x_{i}|\tau_{i},\mathbf{F},\pi_{0})p_{\text{\tiny grid}}(\tau_{1},\dots,\tau_{l}|\rho_{\tau})p_{\text{\tiny rates}}(\mathbf{F}|\mathbf{A},\rho_{f})p(\mathbf{A},\rho_{f})p(\boldsymbol{\pi}_{0}|\rho_{0}).$ (2)
Next, we specify the different components of Eq. 2, starting from the right.
@@
Distribution over initial conditions. The distribution $p(\boldsymbol{\pi}_{0}|\rho_{0})$, with hyperparameter $\rho_{0}$, is defined over the $C$-simplex, and encodes our beliefs about the initial state (i.e. the preparation) of the system. It enters the master equation as the class probabilities of the categorical distribution over the states of the system, at the start of the process. That is $p_{\text{\tiny MJP}}(x,t=0)=\text{Cat}(\boldsymbol{\pi}_{0})$. We either choose $\boldsymbol{\pi}_{0}$ to be the class probabilities of the stationary distribution of the process, or sample it from a Dirichlet distribution. Appendix B provides the specifics.
Distribution over rate matrices. The distribution $p_{\text{\tiny rates}}(\mathbf{F}|\mathbf{A},\rho_{f})$ over the rate matrices encodes our beliefs about the class of MJPs we expect to find in practice. We define it to cover MJPs with state spaces whose sizes range from 2 until $C$, because we want our FIM to be able to handle processes taking values in all those spaces. The distribution is conditioned on the adjacency matrix $\mathbf{A}$, which encodes only connected state spaces (i.e. irreducible embedded Markov chains only), and a hyperparameter $\rho_{f}$ which encodes the range of rate values within the ensemble. Specifically, we define the transition rates as $F_{ij}=a_{ij}f_{ij}$, where $a_{ij}$ is the corresponding entry of $\mathbf{A}$ and $f_{ij}$ is sampled from a set of Beta distributions, with different hyperparameters $\rho_{f}$. Note that these choices restrict the values of the transition rates within the ensemble to the interval $(0,1)$ and hence, they restrict the number of resolvable transitions within the time horizon $T$ of the simulation. We refer the reader to Appendix B, where we specify the prior $p(\mathbf{A},\rho_{f})=p(\mathbf{A})p(\rho_{f})$ and its consequences, as well as give details about the sampling procedure. We also discuss the main limitations of choosing a Beta prior over the transition rates in Section 6.
@@
Distribution over observation grids. The distribution $p_{\text{\tiny grid}}(\tau_{1},\dots,\tau_{l}|\rho_{\tau})$, with hyperparameter $\rho_{\tau}$, gives the probability of observing the MJP at the times $\tau_{1},\dots,\tau_{l}$, and thus encodes our uncertainty about the recording process. Given that we do not know a priori whether the data will be recorded regularly or irregularly in time, nor we know its recording frequency, we define this distribution to cover both regular and irregular cases, as well as various recording frequencies. Note that the number of observation points on the grid is variable. Please see Appendix B for details.
@@
Distribution over noise process. Just as the (instantaneous) solution of the master equation $p_{\text{\tiny MJP}}(x|t,\mathbf{F},\boldsymbol{\pi}_{0})$, the noise distribution $p_{\text{\tiny noise}}(x^{\prime}|x,\rho_{x})$, with hyperparameter $\rho_{x}$, is defined over the set of metastable states $\mathcal{X}$. Recall that FIM solves the MJP inference problem directly in coarse-grained space. The noise distributions then encodes both, possible measurement errors that propagate through the coarse-grained representation, or noise in the coarse-grained representation itself. We provide details of its implementation in Appendix B.
@@
We use the generative model, Eq. 2 above, to generate $N$ MJPs, taking values on state spaces with sizes ranging from 2 to $C$. We then sample $K$ paths per MJP, with probability $p(K)$, on the interval $[0,T]$. The $j$th instance of the dataset thus consists of $K$ paths and is given by
$\mathbf{F}_{j}\sim p_{\text{\tiny rates}}(\mathbf{F}|\mathbf{A}_{j},\rho_{fj}),\,\,\text{and}\,\,\,\boldsymbol{\pi}_{0j}\sim p(\boldsymbol{\pi}_{0}|\rho_{0}),\,\,\text{with}\,\,(\mathbf{A}_{j},\rho_{fj})\sim p(\mathbf{A},\rho_{f}),$
$\text{so that}\quad\Big{\{}X_{jk}(t)\Big{\}}_{k=1}^{K}\sim\text{Gillespie}(\mathbf{F}_{j},\boldsymbol{\pi}_{0j}),$ (3)
@@
3.2 Supervised Recognition Model
In this subsection we introduce a neural recognition model that processes a set of $K$ time series of the form $\{(x_{k1}^{\prime},\tau_{k1}),\dots,(x_{kl}^{\prime},\tau_{kl})\}_{k=1}^{K}$, as generated by the procedure in Eq. 3 above, and estimates the intensity rate matrix $\mathbf{F}$ and initial distribution $\boldsymbol{\pi}_{0}$ of the hidden MJP. Practically speaking, we would like the model to be able to infer MJPs from time series with observation times on any scale. To ensure this, we first normalize all observation times to lie on the unit interval, by dividing them by the maximum observation time $\tau_{\text{\tiny max}}=\text{max}\{\tau_{k1},\dots,\tau_{kl}\}_{k=1}^{K}$, and then rescale the output of the model accordingly (see Appendix C for details).